# Efficient Adversarial Training in LLMs with Continuous Attacks

### v1

@@ A.1 Adversarial Training
The C-AdvUL algorithm has $5$ important hyperparameters, the weight of the utility loss $\alpha_{u}$, toward loss $\alpha_{t}$, and away loss $\alpha_{a}$. Moreover, in preliminary experiments, we observed that away loss tends to dominate the training objective. Models that show very high away loss generally overfitted to the safety objective and stopped answering benign requests. We notice similar issues with the toward loss. Thus, we define a threshold for the away loss $a_{cut}$ and toward loss $t_{cut}$, clamping values below a certain value. If not otherwise defined, we use the following hyperparameters in all experiments. We set $\alpha_{u}=1.0$, $\alpha_{t}=0.5$, and $\alpha_{a}=0.5$, as in [6]. Further, we set $a_{cut}=-5$ and $t_{cut}=0.5$. We use a ratio of $7:1$ for utility and harmful examples during training.
To prevent overfitting in the proposed C-AdvIPO, we use the IPO loss function [31]. Additionally, we set the $\beta$ parameter of IPO to $0.25$ for Gemma models, $0.5$ for Phi-3-Mini, and $X$ for Mistral-7B, which we observed to result in good trade-offs between robustness and utility in preliminary experiments.
@@ Appendix B Robustness extrapolation to discrete attacks
Table 4 summarizes the main adversarial training results. The proposed C-AdvUL and C-AdvIPO algorithms achieve competitive or even superior robustness utility trade-offs compared to the discrete adversarial training algorithm R2D2 [6].
@@ B.1 One-Step Adversarial Training
As a preliminary experiment for scaling continuous adversarial training, we evaluated if C-AdvIPO yields robustness gains if the attack iterations are reduced to one during training. Table 5 illustrates that one-step C-AdvIPO achieves similar robustness improvements as the multi-step variant. Note, that we used the same hyperparameters for the one-step attacks as for the multi-step attack, except for the attack iterations and step size. Further hyperparameter tuning or borrowing recent advances in one-step AT from other domains may help to close this gap [46]. Due to the large computational complexity of attack evaluations, we conduct this experiment on GCG.
@@ Appendix C Adversarial Training Ablations
In C-AdvIPO, the $\beta$ parameter inversely relates to the difference in log-likelihood ratios between the safe answer and the harmful response. Thus, a smaller $\beta$ indicates a larger disparity in these log-likelihood ratios. This intuitively should lead to robustness and utility trade-offs. The left plot in Figure 3 shows the impact of different IPO $\beta$ values on robustness and utility. With $\beta$ values ranging from 00 to $0.5$, a consistent decrease in GCG loss is observed, starting from $6.1$ and dropping to $0.8$. Meanwhile, the MMLU score increases from about $0.25$ to $0.38$. This trend aligns with our expectations and suggests that higher $\beta$ values are associated with lower GCG loss and improved utility, indicating that tuning $\beta$ is crucial for optimizing the robustness-utility trade-off in C-AdvIPO.
@@ Appendix D Adversarial training computational effort
C-AdvUL & C-AdvIPO. The total number of forward passes $F_{UL}$ required by our continuous adversarial training algorithm was calculated as follows.
@@
C-AdvUL Total. The total number of forward passes $F_{UL}$ required by C-AdvUL was calculated as follows.
@@
C-AdvIPO Total. The total number of forward passes $F_{IPO}$ required by C-AdvIPO was calculated as follows.
$F_{IPO}=(2\cdot b_{adv}+b_{adv}\cdot I_{A})\cdot I_{T}.$
@@
$W_{IPO}=(2\cdot b_{adv}+b_{adv}\cdot I_{A})\cdot I_{T}.$

### v2

@@ A.1 Adversarial Training
The CAT algorithm has $5$ important hyperparameters, the weight of the utility loss $\alpha_{u}$, toward loss $\alpha_{t}$, and away loss $\alpha_{a}$. Moreover, in preliminary experiments, we observed that away loss tends to dominate the training objective. Models that show very high away loss generally overfitted to the safety objective and stopped answering benign requests. We notice similar issues with the toward loss. Thus, we define a threshold for the away loss $a_{cut}$ and toward loss $t_{cut}$, clamping values below a certain value. If not otherwise defined, we use the following hyperparameters in all experiments. We set $\alpha_{u}=1.0$, $\alpha_{t}=0.5$, and $\alpha_{a}=0.5$, as in [6]. Further, we set $a_{cut}=-5$ and $t_{cut}=0.5$. We use a ratio of $7:1$ for utility and harmful examples during training.
To prevent overfitting in the proposed CAPO, we use the IPO loss function [31]. Additionally, we set the $\beta$ parameter of IPO to $0.25$ for Gemma models, $0.5$ for Phi-3-Mini, and $X$ for Mistral-7B, which we observed to result in good trade-offs between robustness and utility in preliminary experiments.
@@ Appendix B Robustness extrapolation to discrete attacks
Table 4 summarizes the main adversarial training results. The proposed CAT and CAPO algorithms achieve competitive or even superior robustness utility trade-offs compared to the discrete adversarial training algorithm R2D2 [6].
@@ B.1 One-Step Adversarial Training
As a preliminary experiment for scaling continuous adversarial training, we evaluated if CAPO yields robustness gains if the attack iterations are reduced to one during training. Table 5 illustrates that one-step CAPO achieves similar robustness improvements as the multi-step variant. Note, that we used the same hyperparameters for the one-step attacks as for the multi-step attack, except for the attack iterations and step size. Further hyperparameter tuning or borrowing recent advances in one-step AT from other domains may help to close this gap [46]. Due to the large computational complexity of attack evaluations, we conduct this experiment on GCG.
@@ Appendix C Adversarial Training Ablations
In CAPO, the $\beta$ parameter inversely relates to the difference in log-likelihood ratios between the safe answer and the harmful response. Thus, a smaller $\beta$ indicates a larger disparity in these log-likelihood ratios. This intuitively should lead to robustness and utility trade-offs. The left plot in Figure 3 shows the impact of different IPO $\beta$ values on robustness and utility. With $\beta$ values ranging from 00 to $0.5$, a consistent decrease in GCG loss is observed, starting from $6.1$ and dropping to $0.8$. Meanwhile, the MMLU score increases from about $0.25$ to $0.38$. This trend aligns with our expectations and suggests that higher $\beta$ values are associated with lower GCG loss and improved utility, indicating that tuning $\beta$ is crucial for optimizing the robustness-utility trade-off in CAPO.
@@ Appendix D Adversarial training computational effort
CAT & CAPO. The total number of forward passes $F_{UL}$ required by our continuous adversarial training algorithm was calculated as follows.
@@
CAT Total. The total number of forward passes $F_{UL}$ required by CAT was calculated as follows.
@@
CAPO Total. The total number of forward passes $F_{CAPO}$ required by CAPO was calculated as follows.
$F_{CAPO}=(2\cdot b_{adv}+b_{adv}\cdot I_{A})\cdot I_{T}.$
@@
$W_{CAPO}=(2\cdot b_{adv}+b_{adv}\cdot I_{A})\cdot I_{T}.$